OpenStudy (anonymous):

Simplify completely: 4x^2+5x+3/x

OpenStudy (anonymous):

Is it one whole fraction where (4x^2+5x+3) is the numerator? If it is, then this is how you're meant to simplify: Separate into three whole parts that corresponds to the three terms on the numerator: \[\frac{4x^2+5x+3}{x}=\frac{4x^2}{x}+\frac{5x}{x}+\frac{3}{x}\] Then simplify each fraction: \[\large =4x+5+\frac{3}{x}\]

OpenStudy (anonymous):

how do i simplify a fraction? @Azteck

OpenStudy (anonymous):

When you have a fraction such as this: \[\frac{4x^2}{x}\] It can be rewritten as this right?: \[\frac{4\times x\times x}{x}\]

OpenStudy (anonymous):

so 4x?

OpenStudy (anonymous):

or 16x?

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

I was asking you a yes or no question if you can read what I wrote above. I asked you whether you know that the fraction can be rewritten what I wrote above.

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

So then if I take the 4x out of the numerator and put brackets around the resulting fraction, it can also be rewritten as this right? \[\large \frac{4\times x\times x}{x}=(\frac{4x \times x}{x})\] \[\large =4x(\frac{x}{x})\]

OpenStudy (anonymous):

so what is it simplified?

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

4x^2 + 5x^2 + 3x?

OpenStudy (anonymous):

WHy must you get ahead of yourself? I'm giving you simple steps to follow but you always have to leap ahead, then you hit a wall and ask me whether it's this. Let's just follow a simple road where you understand what you're doing step by step.

OpenStudy (anonymous):

fin @Azteck

OpenStudy (anonymous):

And you don't need to keep tagging my name. I'm already here.

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

Okay, so what I said above, do you agree? Please be honest.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So inside the brackets: \[\large \frac{x}{x}=1\] Correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Then: \[\large 4x(1)=4x\] Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

SO now: \[\large \frac{4x^2}{x}=4x\] Could you tell me what \[\large \frac{5x}{x}\] would equal to?

OpenStudy (anonymous):

alli_bear22, are you there?

OpenStudy (anonymous):

I'm waiting for your response.

OpenStudy (anonymous):

yea srry hold on

OpenStudy (anonymous):

5x?

OpenStudy (anonymous):

Remember: \[\frac{4x^2}{x}=4x\] What would \[\frac{4x}{x}=?\]

OpenStudy (anonymous):

4x or 8x

OpenStudy (anonymous):

you cancel the like-terms. Do you know what like-terms are?

OpenStudy (anonymous):

yea so it would just be 4

OpenStudy (anonymous):

Yes!. So what would \[\frac{5x}{x}=?\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Correct! So now you simplified the question. THat's basically all you need to do.

OpenStudy (anonymous):

\[\frac{3}{x}\] can't be simplified any further, so it's just that.

OpenStudy (anonymous):

so its 4+5+3x?

OpenStudy (anonymous):

but i know im wrong because thats not an option

OpenStudy (anonymous):

Umm... \[\frac{4x^2}{x}+\frac{5x}{x}+\frac{3}{x}=4x+5+\frac{3}{x}\]